OpenStudy (anonymous):

Write the following expression using a single exponent. (5^5)2 ×5^6 5^13 5^16 5^4 5^42

OpenStudy (anonymous):

I think you need to rewrite your problem as \[5^{5} \times 25^{6}\] is not equivalent to any of your answers.

OpenStudy (anonymous):

ahhh ok so your problem is \[(5^{5})^{2} \times 5^{6}\] is that right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok so to figure out how to simplify the (5^5)^2 you need to know this rule: \[(a^{x})^{y} = a^{x\times y}\] so how would you apply this to your problem?

OpenStudy (anonymous):

9765625

OpenStudy (anonymous):

ok...you plugged your equation into a calculator...

OpenStudy (anonymous):

no i got the answer now

OpenStudy (anonymous):

ok good!